OpenStudy (kitkat16):

Help please!!!! A machine needs 2 of its gears replaced. You find 2 gears, one manufactured 5 years ago and the other manufactured 10 years ago. The specifications sheet says that the chance of failure is 5% after 5 years and increases to 8% after 10 years. If the gears on the machine are in no way linked to each other, what are the chances that both will fail? Some one said 0.2% but I can't fiqure out how they got that answer. Steps would be nice.

OpenStudy (kitkat16):

P= 0.05 * 0.08 = 0.004 = 0.4% So they would both fail yes no ???? anyone :(

OpenStudy (kitkat16):

I hate math :(

OpenStudy (kitkat16):

@ganeshie8 check please!!!

OpenStudy (kitkat16):

@Preetha can you check my work please

ganeshie8 (ganeshie8):

If the gears on the machine are in no way linked to each other, what are the chances that both will fail?

ganeshie8 (ganeshie8):

that means the events are "independent"

ganeshie8 (ganeshie8):

for independent events, we have : \(\large P(A \cap B) = P(A) \cdot P(B) \)

ganeshie8 (ganeshie8):

right ?

ganeshie8 (ganeshie8):

Probability for both failing = \(0.05 * 0.08 = 0.004 = 0.4\% \)

OpenStudy (kitkat16):

yes. thank you!!!!

ganeshie8 (ganeshie8):

np :)